1. An autobike begins to on a level road directly toward a building that is 150 feet tall. How far does the
car travel during the time that the angle of elevation from the autobike to the top of the building changes from 27° to 36°
1
Expert's answer
2019-10-07T09:50:41-0400

"d=150cot27^o-150cot36^o=150(1.963-1.376)=88 m."